OpenStudy (anonymous):

Simplify the expression (4^-5(64)^3)^2, expressing the answer as a power of 4 ??????????????????????????????

OpenStudy (mathstudent55):

Is this the expression? \( (4^{-5}(64)^3 )^2 \)

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

Hint: \(64 = 4^3\)

OpenStudy (anonymous):

4^8 power?

OpenStudy (mathstudent55):

Substitute \(4^3\) for \(64\): \( (4^{-5}(4^3)^3)^2 \)

OpenStudy (mathstudent55):

\( (4^{-5} (4^9) )^2 \) \(= 4^{-10} 4^{18} \) \( = 4^8 \) You are correct.

OpenStudy (anonymous):

thanks :)